# A Roadmap for Software Testing in Open Collaborative Development Environments

### v1

@@ 4. Technology Related
In the realm of technology, advancements in software quality assurance are driven by innovative approaches and tools tailored to the demands of open collaborative development environments. From ensuring the integrity of third-party libraries to leveraging crowd intelligence for testing, technological solutions are essential for maintaining software quality amidst the complexities of collaborative development. Through automated identification of vulnerable libraries and harnessing crowd-contributed data for testing enhancements, these technological interventions aim to fortify software ecosystems and enhance overall quality assurance practices. Another significant change has been the rise and prevalence of Artificial Intelligence (AI) algorithms and applications, which has spurred the development of various testing techniques for AI models.
4.1. Quality Assurance for Software Ecosystem
Third-party libraries play a key role in software development because they allow developers to reuse common functionalities instead of reinventing the wheel and substantially improve the productivity of developers. However, outdated third-party libraries can be commonly used but seldom updated, or updated in a delayed way, which might cause software bugs or increase attack surfaces in client projects. Huang et al. (Xi et al., 2019; Huang et al., 2022) conducted large-scale analyses to characterize usages, updates and risks of third-party libraries in Java open-source projects, and proposed a security bug-driven altering system to assist developers for making confident decisions about third-party library version updates. There are studies focusing on automating the identification and updating of deprecated third-party libraries, i.e., APIs. Some of them employed pattern-based techniques, i.e., they rely on the replacement messages in API documents to derive the update patterns, which are subsequently used to guide the replacement of the deprecated APIs (Huang et al., 2021a). Others proposed automating API updates by mining update patterns from the code change history (Fazzini et al., 2019; Haryono et al., 2021). Nevertheless, this source of information only covers a small portion of the deprecated APIs, and is usually delayed.
Software composition analysis is also widely adopted to identify vulnerable libraries used in software applications. For example, Dong et al. (Dong et al., 2019) leveraged named entity recognition to recognize affected libraries from vulnerability description. Chen et al. (Chen et al., 2020) and Lyu et al. (Lyu et al., 2023) formulated this affected library identification problem as an extreme multi-label learning problem, which use library coordinates as labels, and predict labels for each vulnerability based on features extracted from vulnerability description and direct references. Wu et al. (Wu et al., 2024) identified affected libraries and their ecosystems for vulnerabilities via a learning-to-rank technique.
@@ 4.2. Testing Benefited from Crowd Intelligence
In the context of open collaborative software development, a vast amount of data contributed by developers with diverse backgrounds is aggregated. This data encapsulates rich knowledge about software quality assurance, and can be harnessed to empower the testing techniques. For example, Mao et al. (Mao et al., 2017) extracted the test scripts from crowdbased testing to automatically infer the reusable high-level event sequences for enhancing the automated mobile testing. Liu et al. (Liu et al., 2021) collected a large number of GUI screenshots with UI display issues from the crowdsourced testing platform, and used them to train a visual understanding model for automatically detecting the GUIs with display issues and locating the buggy region. Wei et al. (Wei et al., 2022) mined the code/models from open source to obtain the code snippets from the deep learning library documentation, library developer tests, and deep learning models in the wild, then leveraged these information to perform fuzz testing for deep learning libraries.
Figure 2. Distribution of testing tasks with LLMs (Wang et al., 2024b)
@@
In recent years, the pervasive advancements in pre-trained Large Language Models (LLMs) have profoundly impacted various domains, including software testing. These models have been increasingly harnessed to bolster testing capabilities across different facets, ranging from enhancing test coverage in unit testing to diversifying test case generation in integration testing. There is a relevant literature review titled “Software Testing with Large Language Models: Survey, Landscape, and Vision” (Wang et al., 2024b), which provides a comprehensive overview of the utilization of LLMs in software testing. It analyzes 102 relevant studies that have employed LLMs for software testing, examining them from both the software testing and LLMs perspectives. As demonstrated in Figure 2, LLMs are commonly used for test case preparation (including unit test case generation, test oracle generation, and system test input generation), program debugging, and bug repair, while there is currently no practices for applying LLMs in the tasks of early testing life-cycle (such as test requirement, test plan, etc).
@@
For unit test case generation, a majority of the earlier published studies adopt the pre-training or fine-tuning schema, for example Alagarsamy et al. (Alagarsamy et al., 2023) first pre-trained the LLM with the focal method and asserted statements to enable the LLM to have a stronger foundation knowledge of assertions, then fine-tuned the LLM for the test case generation task where the objective is to learn the relationship between the focal method and the corresponding test case. By comparison, most later studies typically focus on how to design the prompt, to make the LLM better understanding the context of the task. Yuan et al. (Yuan et al., 2023) performed an empirical study to evaluate ChatGPT’s capability of unit test generation with both a quantitative analysis and a user study in terms of correctness, sufficiency, readability, and usability. And results show that the generated tests still suffer from correctness issues, including diverse compilation errors and execution failures. They then proposed an approach that leveraged the ChatGPT itself to improve the quality of its generated tests with an initial test generator and an iterative test refiner. Specifically, the iterative test refiner iteratively fixed the compilation errors in the tests generated by the initial test generator, which follows a validate-and-fix paradigm to prompt the LLM based on the compilation error messages and additional code context.
@@
For system test input generation, it varies for specific types of software being tested. For example, for mobile applications, the test input generation requires providing a diverse range of text inputs or operation combinations (e.g., click a button, long press a list) (Liu et al., 2023a, 2024), which is the key to testing the application’s functionality and user interface; while for Deep Learning (DL) libraries, the test input is a program which covers diversified DL APIs (Deng et al., 2023b, a). For example, Liu et al. (Liu et al., 2024) formulates the test input generation of mobile GUI testing problem as a Q&A task, which asks LLM to chat with the mobile apps by passing the GUI page information to LLM to elicit testing scripts (i.e., GUI operation), and executing them to keep passing the app feedback to LLM, iterating the whole process. The proposed GPTDroid also introduces a functionality-aware memory prompting mechanism that equips the LLM with the ability to retain testing knowledge of the whole process and conduct long-term functionality-based reasoning to guide exploration. Deng et al. (Deng et al., 2023b) used both generative and infilling LLMs to generate and mutate valid/diverse input DL programs for fuzzing DL libraries. In detail, it first uses a generative LLM (CodeX) to generate a set of seed programs (i.e., code snippets that use the target DL APIs). Then it replaces part of the seed program with masked tokens using different mutation operators and leverages the ability of infilling LLM (InCoder) to perform code infilling to generate new code that replaces the masked tokens.
@@ 4.3. Testing for AI Models and Applications
The rise of Artificial Intelligence (AI) applications raises concerns about trustworthiness, particularly in safety-critical domains such as self-driving systems and medical treatments. Software testing plays a crucial role in detecting and addressing discrepancies between expected and actual behaviors in these applications. However, testing AI systems presents unique challenges due to their statistical nature, evolving behavior, and the oracle problem.
@@
With the advancement of LLMs, there are also studies focusing on the robustness evaluation of LLMs. For example, Wan et. al. (Wan et al., 2024) developed an approach for triggering logical reasoning failures in LLMs. Wang et. al. (Wang et al., 2024a) proposed to automatically detect the issues related with social bias in image generation models.
@@ 5. Challenges and Opportunities
Despite the advancements and breakthroughs discussed in the preceding sections, open collaborative development environments continue to present numerous challenges, which suggests ongoing research opportunities and emerging trends in the coming years.
@@ 5.1. Leveraging LLMs for Enhancing Testing
The first is the challenge for achieving high coverage. Exploring the diverse behaviors of the software under test to achieve high coverage is always a significant concern in software testing. In this context, test generation differs from code generation, as code generation primarily focuses on producing a single, correct code snippet, whereas software testing requires generating diverse test inputs to ensure better coverage of the software. Although setting a high temperature can facilitate the LLMs in generating different outputs, it remains challenging for LLMs to directly achieve the required diversity. For example, for unit test case generation, in SF110 dataset, the line coverage is merely 2% (Siddiq et al., 2023). For system test input generation, in terms of fuzzing DL libraries, the API coverage for TensorFlow is reported to be 66% (2215/3316) (Deng et al., 2023b). Current studies typically utilize mutation testing together with the LLMs to generate more diversified outputs. For example, when fuzzing a DL library, instead of directly generating the code snippet with LLM, Deng et al. (Deng et al., 2023b) replace parts of the selected seed (code generated by LLM) with masked tokens using different mutation operators to produce masked inputs. They then leverage the LLM to perform code infilling to generate new code that replaces the masked tokens, which can significantly increase the diversity of the generated tests. Other potential research direction could involve utilizing testing-specific data to train or fine-tune a specialized LLM that is specifically designed to understand the nature of testing. By doing so, the LLM can inherently acknowledge the requirements of testing and autonomously generate diverse outputs.
@@
The second is the challenge in test oracle problem. The oracle problem has been a longstanding challenge in various testing applications, e.g., testing machine learning systems (Zhang et al., 2022b) and testing deep learning libraries (Deng et al., 2023b). To alleviate the oracle problem to the overall testing activities, a common practice is to transform it into a more easily derived form, often by utilizing differential testing (Sun et al., 2023) or focusing on only identifying crash bugs (Liu et al., 2024). Exploring the use of LLMs to derive other types of test oracles represents an interesting and valuable research direction. Specifically, metamorphic testing is also widely used in software testing practices to help mitigate the oracle problem, yet in most cases, defining metamorphic relations relies on human ingenuity. Another promising avenue is exploring the capability of LLMs to automatically generate test cases based on metamorphic relations, covering a wide range of inputs. The advancement of multi-model LLMs like GPT-4 may open up possibilities for exploring their ability to detect bugs in software user interfaces and assist in deriving test oracles. By leveraging the image understanding and reasoning capabilities of these models, one can investigate their potential to automatically identify inconsistencies, errors, or usability issues in user interfaces.
@@
The third is the challenge in the real-world application of LLMs in software testing. Due to concerns regarding data privacy, when considering real-world practice, most software organizations tend to avoid using commercial LLMs and would prefer to adopt open-source ones with training or fine-tuning using organization-specific data. Furthermore, some companies also consider the current limitations in terms of computational power or pay close attention to energy consumption, they tend to fine-tune medium-sized models. It might be quite challenging for these models to achieve similar performance to what existing papers have reported. Recent research has highlighted the importance of high-quality training data in improving the performance of models for code-related tasks [152], yet manually building high-quality organization-specific datasets for training or fine-tuning is time-consuming and labor-intensive. To address this, one is encouraged to utilize the automated techniques of mining software repositories to build the datasets, for example, techniques like key information extraction techniques from Stack Overflow (Shi et al., 2019) offer potential solutions for automatically gathering relevant data. In addition, exploring the methodology for better fine-tuning the LLMs with software-specific data is worth considering because software-specific data differs from natural language data as it contains more structural information, such as data flow and control flow. Previous research on code representations has shown the benefits of incorporating data flow (Guo et al., 2021).
5.2. Testing for LLMs and Related Applications
There have been numerous research efforts on testing machine learning and deep learning models, however, in the context of general artificial intelligence, i.e., LLMs, there has been relatively less exploration. Example researches include robustness evaluation of LLMs in logical reasoning tasks (Wan et al., 2024), evaluations of LLMs in task automation (Shen et al., 2023), heterogeneous value alignment evaluation of LLMs (Zhang et al., 2023b), etc. However, there are still many other dimensions of capabilities in LLMs that need to be tested and evaluated, such as contextual understanding and situational comprehension. Additionally, there is a lack of systematic testing methodology specifically designed for LLMs. Furthermore, with the emergence of multi-modal LLMs, there is a need for more research and attention on testing such models.
Furthermore, LLMs are not only utilized in isolation but are also integrated into applications, with a typical example being the LLM-as-agent paradigm (Xi et al., 2023; Liu et al., 2023b). In this scenario, LLMs engage in cognitive understanding, decision-making, and other tasks to support specific application scenarios. This integration extends beyond just the LLMs itself, and encompasses various external factors. For instance, it involves external environments, memory modules for storing historical information from interactions, external knowledge bases for retrieving the latest information, and tools or external services for utilization. Ensuring the robustness and reliability of LLM-as-agent systems requires testing not only the individual components but also their integration and interaction. This involves verifying the model’s ability to effectively leverage external resources, adapt to dynamic environments, and maintain coherence across diverse tasks and contexts. Additionally, testing strategies should encompass scenarios that stress the system’s capabilities, such as handling unexpected inputs, responding to novel situations, and gracefully recovering from errors or disruptions. By adopting a holistic testing approach, developers can gain confidence in the functionality and performance of LLM-as-agent systems across various real-world scenarios.
GPTs, as a new form of service based on LLMs, will make the GPT Store a new channel for people to access applications, similar to the Google Play Store or the Apple App Store in the era of mobile applications. They are custom versions of GPT tailored for specific purposes, allowing users to create personalized iterations of GPT to better suit their needs (OpenAI, 2024). These customized GPTs can be designed for various tasks, such as teaching children math, providing assistance in board games, or generating stickers. Therefore, ensuring the quality and reliability of all GPTs on the platform is a critical concern (Zhang et al., 2024). For example, there may be instances where certain GPTs engage in illegal activities such as unauthorized access to user privacy. Effective testing mechanisms must be established to address such concerns. Moreover, discrepancies between the descriptions of GPT capabilities and their actual performance may arise, necessitating robust testing procedures to ensure consistency and accuracy. Additionally, leveraging user-contributed feedback and reviews on the GPT Store platform could serve as a valuable resource for identifying defects and improving GPT performance.
5.3. Quality Assurance for Auto-generated Code

### v2

@@ 4. Technology Related
Advancements in AI technology have significantly enhanced testing capabilities, introducing intelligent automation and improving efficiency in open-collaborative environments. At the same time, the rise of AI-driven applications and systems has created new demands for testing, necessitating specialized techniques to validate the functionality, robustness, and fairness of AI models and applications. These dual developments highlight the evolving role of technology in both strengthening traditional testing processes and addressing the unique challenges posed by AI-centric systems.
@@ 4.1. Testing Benefited from Crowd Intelligence
In the context of open collaborative software development, a vast amount of data contributed by developers with diverse backgrounds is aggregated. This data encapsulates rich knowledge about software quality assurance, and can be harnessed to empower the testing techniques. For example, Mao et al. (Mao et al., 2017) extracted the test scripts from crowdbased testing to automatically infer the reusable high-level event sequences for enhancing the automated mobile testing. Liu et al. (Liu et al., 2021) collected a large number of GUI screenshots with UI display issues from the crowdsourced testing platform, and used them to train a visual understanding model for automatically detecting the GUIs with display issues and locating the buggy region. Wei et al. (Wei et al., 2022) mined the code/models from open source to obtain the code snippets from the deep learning library documentation, library developer tests, and deep learning models in the wild, then leveraged this information to perform fuzz testing for deep learning libraries.
4.2. Testing with AI Models
Figure 2. Distribution of testing tasks with LLMs (Wang et al., 2024a)
Since the introduction of the concept of training “Deep Neural Network” by Geoffrey Hinton in 2006, deep learning has been increasingly adopted to develop cutting-edge tools for software testing research, thanks to its ability to enhance performance. A comprehensive survey (Yang et al., 2022) explores how deep learning is applied in testing areas such as test case generation, bug localization, and application testing.
@@ 4.2. Testing with AI Models
In recent years, the pervasive advancements in LLMs have profoundly impacted various domains, including software testing. These models have been increasingly harnessed to bolster testing capabilities across different facets, ranging from enhancing test coverage in unit testing to diversifying test case generation in integration testing. There is a relevant literature review titled “Software Testing with Large Language Models: Survey, Landscape, and Vision” (Wang et al., 2024a), which provides a comprehensive overview of the utilization of LLMs in software testing. It analyzes 102 relevant studies that have employed LLMs for software testing, examining them from both the software testing and LLMs perspectives. As demonstrated in Figure 2, LLMs are commonly used for test case preparation (including unit test case generation, test oracle generation, and system test input generation), program debugging, and bug repair. However, there is currently no practices for applying LLMs in the tasks of early testing life-cycle (such as test requirement, test plan, etc).
@@
For unit test case generation, a majority of the earlier published studies adopt the pre-training or fine-tuning schema. For example Alagarsamy et al. (Alagarsamy et al., 2023) first pre-trained the LLM with the focal method and asserted statements to enable the LLM to have a stronger foundation knowledge of assertions. They then fine-tuned the LLM for the test case generation task where the objective is to learn the relationship between the focal method and the corresponding test case. By comparison, most later studies typically focus on how to design the prompt, to make the LLM better understanding the context of the task. Yuan et al. (Yuan et al., 2023) performed an empirical study to evaluate ChatGPT’s capability of unit test generation with both a quantitative analysis and a user study in terms of correctness, sufficiency, readability, and usability. And results show that the generated tests still suffer from correctness issues, including diverse compilation errors and execution failures. To address this, they proposed an approach where ChatGPT was used to improve the quality of its own generated tests, using an initial test generator and an iterative test refiner. The iterative refiner followed a validate-and-fix approach, correcting compilation errors by prompting the LLM based on error messages and additional code context.
@@
For system test input generation, it varies for specific types of software being tested. For mobile applications, test input generation requires a wide range of text inputs or operation combinations (e.g., clicking a button or long-pressing a list) to test the application’s functionality and user interface (Liu et al., 2023a, 2024a). In contrast, for Deep Learning (DL) libraries, the test input consists of programs that cover diverse DL APIs (Deng et al., 2023b, a). For example, Liu et al. (Liu et al., 2024a) formulates the test input generation of mobile GUI testing problem as a Q&A task, which asks LLM to chat with the mobile apps by passing the GUI page information to the LLM. The LLM then generates testing scripts (i.e., GUI operations) and executes them while receiving app feedback, iterating the process. The proposed GPTDroid also introduces a functionality-aware memory prompting mechanism that equips the LLM with the ability to retain testing knowledge of the whole process and conduct long-term functionality-based reasoning to guide exploration. Deng et al. (Deng et al., 2023b) used both generative and infilling LLMs to generate and mutate valid/diverse input DL programs for fuzzing DL libraries. The process starts with a generative LLM (CodeX) to generate a set of seed programs using target DL APIs. Next, part of the seed program is replaced with masked tokens, and an infilling LLM (InCoder) is used to fill in the masked tokens and generate new code.
@@ 4.3. Testing for AI Models and Applications
The rise of AI applications raises concerns about trustworthiness, particularly in safety-critical domains such as self-driving systems and medical treatments. Software testing plays a crucial role in detecting and addressing discrepancies between expected and actual behaviors in these applications. However, testing AI systems presents unique challenges due to their statistical nature, evolving behavior, and the oracle problem.
@@ 5. Challenges and Opportunities
Despite the advancements and breakthroughs discussed in the preceding sections, open collaborative development environments and the rapid evolution of AI technologies/ecosystems continue to present numerous challenges, highlighting ongoing research opportunities and emerging trends in the coming years.
@@ 5.1. Leveraging LLMs for Enhancing Testing
Tese case generation. Exploring diverse software behaviors while minimizing costs remains a critical challenge in software testing. Test case generation, in particular, poses significant obstacles in traditional automated testing due to the difficulty of producing test inputs with correct grammar and semantics, especially for complex or composite inputs. As a result, automated testing often struggles to achieve comprehensive coverage. LLMs offer promising potential with their exceptional ability to understand software context, yet current research has barely scratched the surface of their capabilities. For example, for unit test case generation, in SF110 dataset, the line coverage is merely 2% (Siddiq et al., 2023). For system test input generation, in terms of fuzzing DL libraries, the API coverage for TensorFlow is reported to be 66% (2215/3316) (Deng et al., 2023b).
As the initial wave of enthusiasm for LLMs has subsided, people have become increasingly aware of the limitations of LLMs and the strengths of program analysis and traditional testing techniques. Consequently, the integration of LLMs with traditional approaches has emerged as an important direction for future exploration. One direction may utilize mutation testing together with the LLMs to generate more diversified outputs. For example, when fuzzing a DL library, instead of directly generating the code snippet with LLM, Deng et al. (Deng et al., 2023b) replace parts of the selected seed (code generated by LLM) with masked tokens using different mutation operators to produce masked inputs. They then leverage the LLM to perform code infilling to generate new code that replaces the masked tokens, which can significantly increase the diversity of the generated tests. Automatically generate test cases based on metamorphic relations to cover a wide range of inputs is also a promise avenue. Combine LLM with other traditional technique is also a promising direction, e.g., Jiang et. al. (Jiang et al., 2024) conduct a systematic study on the LLMs and constraint-based tools for test input generation, and find that there are limitations for LLMs in specific scenarios such as sequential calculation, where constraint-based tools are in a position of strength. By combining them together, there can be 1.4x to 2.3x improvement than the baselines.
Other potential research direction could involve utilizing testing-specific data to train or fine-tune a specialized LLM that is specifically designed to understand the nature of testing. By doing so, the LLM can inherently acknowledge the requirements of testing and autonomously generate diverse outputs.
Test oracle problem. The oracle problem has been a longstanding challenge in various testing applications, e.g., testing machine learning systems (Zhang et al., 2022b) and testing deep learning libraries (Deng et al., 2023b). To alleviate the oracle problem to the overall testing activities, a common practice is to transform it into a more easily derived form, often by utilizing differential testing (Sun et al., 2023) or focusing on only identifying crash bugs (Liu et al., 2024a).
@@
Exploring the use of LLMs to derive other types of test oracles represents an interesting and valuable research direction. Specifically, metamorphic testing is also widely used in software testing practices to help mitigate the oracle problem, yet in most cases, defining metamorphic relations relies on human ingenuity. Recent preliminary attempts have demonstrated the potential of leveraging LLMs to automatically discover and define these relations, offering a promising avenue for further investigation and application (Luu et al., 2023). Another promising avenue is exploring the capability of LLMs to automatically generate test cases based on metamorphic relations, covering a wide range of inputs. Apart from that, the advancement of multi-model LLMs like GPT-4 may open up possibilities for exploring their ability to detect bugs in software user interfaces and assist in deriving test oracles. By leveraging the image understanding and reasoning capabilities of these models, one can investigate their potential to automatically identify inconsistencies, errors, or usability issues in user interfaces, e.g., VisionDroid (Liu et al., 2024b) makes the first attempt towards this direction.
@@
Real-world application of LLMs in software testing. Due to concerns regarding data privacy, when considering real-world practice, most software organizations tend to avoid using commercial LLMs and would prefer to adopt open-source ones with training or fine-tuning using organization-specific data. Furthermore, some companies also consider the current limitations in terms of computational power or pay close attention to energy consumption, they tend to fine-tune medium-sized models. It might be quite challenging for these models to achieve similar performance to what existing papers have reported. Recent research has highlighted the importance of high-quality training data in improving the performance of models for code-related tasks. A notable example in this area is the StarCoder2 model (Lozhkov et al., 2024), where the StarCoder2-15B significantly outperforms other models of similar size (e.g., CodeLlama-13B) and even matches or surpasses the performance of CodeLlama-34B. The superior performance of StarCoder2 is reported as attributing to the curation of high-quality open data sources, such as GitHub issues, pull requests, Kaggle datasets, Jupyter notebooks, and code documentation. Additionally, rigorous data preprocessing steps, including deduplication and the application of filters to eliminate low-quality code, have played a crucial role in enhancing the model’s effectiveness. However, building high-quality, organization-specific datasets for training or fine-tuning is a time-consuming and labor-intensive process. To address this challenge, automated techniques from the field of mining software repositories (Hassan, 2008)—an area that has made significant strides over the past decade—can be employed. These techniques enable efficient extraction and analysis of key information, streamlining the dataset creation process.
In addition, exploring methodologies for better pre-training or fine-tuning LLMs with software-specific data is a promising direction. In recent times, the approach to LLM pre-training and fine-tuning has largely followed the “more data, better results” philosophy, with an emphasis on utilizing as much data as possible. However, the availability of open data is limited, and this path has nearly reached its limits. On the other hand, software-specific data differs from natural language data in that it contains more structural information, such as data flow and control flow. Therefore, enabling LLMs to better learn these software-specific structural elements may become a key focus in future research. Previous work on code representations has highlighted the benefits of incorporating data flow information, as demonstrated by Guo et al. [2021] in their GraphCodeBERT model.
5.2. Testing for LLMs and LLM-centric Applications
Since the emergence of LLMs, AI-enabled software applications have rapidly advanced. The quality of these applications now relies not only on functional correctness but also on the performance of the embedded or associated AI models. This shift has introduced unprecedented challenges to testing.
Testing methodology specifically designed for LLMs. There have been numerous research efforts on testing machine learning and deep learning models, however, in the context of general artificial intelligence, i.e., LLMs, there has been relatively less exploration in software-related conferences and journals. At AI-related conferences, much work has been done on benchmarking LLMs, such as evaluating their performance in task automation (Wang et al., 2024b), instruction tuning (Shen et al., 2024), and judge assistants (Chen et al., 2024). These benchmarking efforts have significantly advanced the field, with some works already having hundreds of citations. From the perspective of software engineering, a systematic testing methodology specifically designed for LLMs is urgently needed. It’s essential to extend testing beyond basic benchmarks and include functional, non-functional, and safety testing, especially considering the complex and unpredictable behaviors of LLMs when applied in real-world software applications. Furthermore, with the emergence of multi-modal LLMs, there is a need for more research and attention on testing such models.
Traditional software testing techniques face several challenges due to the non-deterministic nature, complex input structures, and the lack of transparency in LLMs’ decision-making processes. While techniques like fuzz testing and metamorphic testing may be adapted and extended to these scenarios, they still do not fully address the core complexities of LLMs. These methods may help detect specific issues but often fail to tackle the underlying dynamics of model behavior, such as how models handle ambiguous or unseen inputs. Therefore, there is room for more novel approaches and ideas in testing LLMs, potentially requiring entirely new paradigms that go beyond traditional methods to better evaluate their performance and reliability in diverse, real-world tasks.
Testing LLM-as-agent systems. Whether for classification tasks or generation task, LLMs remain far from mature. Yet, LLM-powered applications have already proliferated, with the LLM-as-agent paradigm (Xi et al., 2023; Liu et al., 2023b) being a prominent example. In this scenario, LLMs handle tasks such as cognitive understanding and decision-making to support specific applications. This paradigm extends beyond the LLMs themselves, incorporating external components like environments, memory modules for storing interaction history, external knowledge bases for retrieving up-to-date information, and tools or services for task execution. Testing the performance of LLM-as-agent systems requires evaluating both traditional software metrics (e.g., response speed, fault tolerance, and reliability) and AI-specific aspects (e.g., compliance, robustness, generalization, trustworthiness, and fairness).
Apart from that, with recent advancements, such as Claude 3.5’s computer use and AutoGLM, LLM agents are demonstrating greater autonomy, more sophisticated tool usage, and emerging visual capabilities, evolving toward LLM-centered operating systems. Testing in this context involves verifying the model’s ability to effectively utilize external resources, adapt to dynamic environments, and maintain coherence across complex tasks. Additionally, testing strategies must account for challenging scenarios, including handling unexpected inputs, adapting to novel situations, and recovering gracefully from errors or disruptions. By adopting a holistic testing approach, developers can ensure the functionality and reliability of LLM-as-agent systems in diverse real-world scenarios.
[…]
5.3. Testing and Quality Assurance for Auto-generated Code